OpenStudy (anonymous):

The volume is 200cm^3 I am trying to work out the change in the height of a cylinder the volume of the cylinder V = pi*r^2*h the rate of change is 1/1000πcm/sec the current radius is 45cm radius is 45cm and the height is increasing at a rate of 1/1000πcm/sec. I need to calculate the increase in size of the radius hence r

OpenStudy (anonymous):

rate of change of what? is it \(r\) that is changing?

OpenStudy (badhi):

I guess the volume doesn't change wrt time. If so, $$V=\pi r^2 h$$ differentiating wrt time - t $$\frac{dV}{dt}=\pi \left(2r\frac{dr}{dt}\right)h+\pi r^2\frac{dh}{dt}\\ 0=\pi r\left(2\frac{dr}{dt}h+r\frac{dh}{dt}\right)\\ \frac{dr}{dt}=-\frac{r}{2h}\frac{dh}{dt}$$

OpenStudy (anonymous):

i updated my answer. Also when i add the values in I guess r = 45 and the dh/dt = 1/1000pi. But dont understand what 2h is. Could someone please point me in the right direction

OpenStudy (badhi):

Ok, h is the height of the cylinder when the radius is r. In the question the volume is given thus, $$V=\pi r^2 h \implies h=\frac{V}{\pi r^2}$$ substituting with the previous derivation, $$\frac{dr}{dt}=-\frac{2r}{\frac{V}{\pi r^2}}\frac{dh}{dt}\\ \frac{dr}{dt}=-\frac{2\pi r^3}{V}\frac{dh}{dt}$$ now all the parameters are provided thus you can find the rate of change in radius

OpenStudy (anonymous):

thanks I understand what todo now